# Privacy-Preserving Dataset Combination

### v1

@@ Abstract.
Access to diverse, high-quality datasets is crucial for machine learning model performance, yet data sharing remains limited by privacy concerns and competitive interests, particularly in regulated domains like healthcare. This dynamic especially disadvantages smaller organizations that lack resources to purchase data or negotiate favorable sharing agreements. We present SecureKL, a privacy-preserving framework that enables organizations to identify beneficial data partnerships without exposing sensitive information. Building on recent advances in dataset combination methods, we develop a secure multiparty computation protocol that maintains strong privacy guarantees while achieving ¿90% correlation with plaintext evaluations. In experiments with real-world hospital data, SecureKL successfully identifies beneficial data partnerships that improve model performance for intensive care unit mortality prediction while preserving data privacy. Our framework provides a practical solution for organizations seeking to leverage collective data resources while maintaining privacy and competitive advantages. These results demonstrate the potential for privacy-preserving data collaboration to advance machine learning applications in high-stakes domains while promoting more equitable access to data resources. Our code is publicly available at https://anonymous.4open.science/r/Private-Preserving_Data_Combination-451E.
isbn: 978-1-4503-XXXX-X/18/06
1. Introduction
@@
Empirical scaling laws have established clear relationships between model performance and three key factors: compute, data, and model size (Hestness2017DeepLS, ; Kaplan2020ScalingLF, ). These relationships have driven remarkable improvements across computer vision, language processing, speech recognition, reinforcement learning, and healthcare applications (sun2017revisiting, ; mahajan2018exploring, ; brown2020language, ; hoffmann2022training, ; park2019specaugment, ; zhang2020pushing, ; ashvin2020accelerating, ; reed2022generalist, ; sheller2020federated, ; mckinney2020international, ). Beyond simple scaling, the diversity of training data has proven crucial for enhancing model robustness to distribution shifts and mitigating performance disparities across demographic groups (shen2024data, ; miller2021accuracy, ).
@@
However, access to data varies significantly across entities and domains. While large tech companies have the data and compute resources to train the foundational models that now dominate general tasks, smaller players often lack such access. Domain-specific data is also becoming increasingly valuable for fine-tuning these general models (alsentzer2019publicly, ; gururangan2020don, ; lee2020biobert, ), creating a competitive advantage for the data owners. As a result, entities with domain-specific data are more reluctant to share it for free, opting instead to sell it in emerging data markets (acemoglu2022too, ; huang2021toward, ; liang2018survey, ). This dynamic particularly disadvantages smaller organizations, which often lack both the resources to purchase data and the leverage to negotiate favorable sharing agreements.
@@
Scaling datasets in regulated domains is particularly challenging due to legal constraints and unpredictable outcomes from altering the training data composition. In healthcare, for example, patient data is heavily regulated by laws such as the Health Insurance Portability and Accountability Act (HIPAA), which imposes strict data-sharing constraints to protect patient privacy. Moreover, accumulating data from multiple sources introduces the risk of domain shift, where data from different distributions may degrade model performance instead of improving it (koh2021wilds, ; wang2022generalizing, ; taori2020measuring, ). Contrary to the intuition that more data always leads to better performance, combining datasets does not guarantee improvements — in fact, performance can decrease when incorporating additional data sources (bradley2021unrepresentative, ; meng2018statistical, ; shen2024data, ). This non-monotonic behavior means that carefully selecting which datasets to combine is crucial, as using all available data may actually perform worse than using an optimal subset.
@@
While existing methods aim to identify datasets combinations that can improve performance, they often assume access to all relevant datasets (shen2024data, ; ilyas2022datamodels, ). However, this is impractical because entities are reluctant to share their data due to privacy risks or the competitive value they associate with it. This reluctance to share creates a bottleneck in improving model performance. In high-stakes settings like healthcare, this bottleneck is particularly detrimental, as access to diverse data has the potential to drive significant advancements in patient care and outcomes (beam2018big, ; ghassemi2019practical, ; chen2020treating, ; beaulieu2019trends, ).
@@
  2. (2)
    We extend KL-XY score (shen2024data, ) with a secure multiparty protocol that enables privacy-preserving evaluation of potential data partnerships. Our protocol maintains strong privacy guarantees while utilizing the complete underlying datasets, achieving ¿90% correlation with plaintext computations.
  3. (3)
@@
  4. (4)
    In simulated scenarios requiring selection of three partner hospitals, Private-KL-XY outperforms alternative selection strategies including demographic-based selection (using gender, race, and age), and limited-sample plaintext selection.
@@ 2. Problem Setup
(Baseline Performance) $\mathrm{AUC}_{o}=\max_{f(\theta)}\,\mathrm{AUC}(\mathcal{M},\mathcal{D}_{o})$
@@
(Combined Performance) $\mathrm{AUC}_{T}=\max_{f(\theta)}\,\mathrm{AUC}(\mathcal{M},\mathcal{D}_{T}).$
@@
(Ideal Dataset Combination) $\pi^{*}(\mathcal{D}_{o},n)=\operatorname*{arg\,max}_{T\subset{[1...N]\choose n}}\mathrm{AUC}_{T}$
@@
Our approach uses Kullback-Leibler (KL)-divergence-based methods to gauge data utility, building on prior work (shen2024data, ). KL divergence (kullback1951information, ), also called information gain (quinlan1986induction, ), describes a measure of how much a model probability distribution $Q$ is different from a true probability distribution $P$:
(Kullback–Leibler Divergence) $\mathrm{KL}(P||Q)=\int_{x\in\mathcal{X}}\log\frac{P(\mathop{}\!\mathrm{d}x)}{Q(\mathop{}\!\mathrm{d}x)}P(\mathop{}\!\mathrm{d}x)$
Because computing KL-divergence on datasets $\mathcal{D}_{o}$ and $\mathcal{D}_{i}$ is non-trivial,  (shen2024data, ) proposes two groups of scores to make this divergence approximation tractable from small samples. Specifically, score $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ first trains a logistic regression model on $\mathcal{D}_{o}\cup\mathcal{D}_{i}$ – where the labels are folded into the covariates — with the goal of inferring dataset membership. Then, the resulting model’s probability score function $\text{Score}(\cdot):\mathcal{X,Y}\to[0,1]$ is averaged over a dataset in $H_{o}$, obtaining
(KL-XY Score) $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}=\mathbb{E}_{(x,y)\sim\mathcal{D}_{o}}(\text{Score}(x,y)).$
@@
We operate under a semi-honest privacy model—also known as honest-but-curious or passive security—where parties follow protocols but may probe intermediate values. Parties are “curious”, meaning that they can probe into the intermediate values to avoid paying for the data. This assumes a weaker security model than malicious security where a corrupted party may input foul data, but ensures the algorithm to be private throughout the computation. This privacy preservation model incentivizes collaboration, improving upon methods in  (shen2024data, ) .
@@
To secure this divergence computation cryptographically, Secure Multiparty Computation (MPC) (yao1982protocols, ; shamir1979share, ) protocols are leveraged. Specifically, in $\mathrm{SecureKL}$, each party encodes $\mathcal{D}_{o}$ and $\mathcal{D}_{i}$ to preserve privacy for both parties. This is implemented with the research framework CrypTen (knott2021crypten, ), specialized for MPC and machine learning. Our algorithmic and engineering details are in Sections  3 and  4, respectively. For related secure techniques, see Section 5.4.
@@
3. Methods
Figure 1. Our Method $\mathrm{SecureKL}(\mathcal{D}_{o},\mathcal{D}_{i})$. Each side encrypts their data, and a model is trained on their joint data. Using private KL-based measures, their distance is computed, and the final result is revealed after decryption, which requires both parties.
[…]
Data is encrypted on each side.
@@ 3. Methods
Category 1, medium-to-high leakage, sharing raw data. $\pi_{s}(n,k)$ supposes each hospital to share a dataset of size $k$; a default setting of $1\%$ is commonplace practice in some contracts, as a pre-requisite to being considered (lds, ). Though leakage can be controlled through $k$, the data is inherently sensitive. The underlying distance uses  (shen2024data, ) ’s KL-XY Score.
@@
3.1. Trivial Baseline: Blind Selection
Blind selection refers to the process when no information is provided. $\pi_{0}(n)$ randomly selects $n$ disjoint data sources, until data purchasing budget runs out. This random strategy may evade selection biases and help gather diverse data. Yet, prior work  (shen2024data, ) suggests that $\pi_{0}(1)$ – randomly selecting one source – for ICU is risky and inefficient for mortality prediction.
3.2. Sharing Summary Statistics
@@
(Demographic-based Strategy) $\pi_{d}(n=1)=\operatorname*{arg\,min}_{i\in[1..N]}\,L_{2}(\phi(\mathcal{D}_{o})||\phi(\mathcal{D}_{i})).$
3.3. KL-based Methods, in Plaintext
@@
This is implemented with KL-based measures similar to (shen2024data, ) . To recap, $\mathrm{KL}(P||Q)$ is not symmetrical, meaning that it is not a ”metric” that satisfies triangle inequality. Intuitively, this means the measure is directional: a hospital’s distribution $P_{o}$ may be ”close” to the target distribution $P_{i}$, but not the other way around:
(Ideal Estimator) $\mathrm{KL}(P_{o}||P_{i})=\int_{x\in\mathcal{X}}\log\frac{P_{o}(\mathop{}\!\mathrm{d}x)}{P_{i}(\mathop{}\!\mathrm{d}x)}P_{i}(\mathop{}\!\mathrm{d}x)$
@@
Specifically, $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ score used in $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ first trains a logistic regression model (Cox1958TheRA, ) on $\mathcal{D}_{o}\cup\mathcal{D}_{i}$ – where the labels are folded into the covaraiates – with the goal of inferring dataset membership. A score of $0.5$ or less means the datasets are not distinguishable, making the data potentially useful.  (shen2024data, ) established the insight that in data-limited domains of heterogeneous data sources, domain shifts of the covariates are useful for predicting whether the additional data helps the original task. We note again that even though this model is trained on both parties’ data, the final algorithm that the hospital uses to train on combined data is not restricted.
@@
(KL-X) $\mathrm{KL}_{\mathcal{X}}=\mathbb{E}_{(x)\sim\mathcal{D}_{o}}(\text{Score}(x)).$
(KL-XY) $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}=\mathbb{E}_{(x,y)\sim\mathcal{D}_{o}}(\text{Score}(x,y)).$
We focus on $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$, and reproduce (shen2024data, ) ’s results that it is predictive of downstream change in AUC.
@@
(KL-based Strategy, in plaintext) $\pi_{s}(n=1,k=K)=\operatorname*{arg\,min}_{i\in[1..N]}\,g_{\mathrm{KL}}(\mathcal{D}_{o},\mathcal{D}_{i}).$
@@
3.4. SecureKL: Private KL-based Method
Using MPC, we extend on $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ to require no information sharing (besides what was already assumed public). Specifically we leverage the MPC based framework provided by CrypTen (knott2021crypten, ), a library designed for privacy-preserving machine learning, to implement private $\mathrm{KL}_{\mathcal{X}}$ and $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$. As illustrated in Figure 1, the logistic regression as well as the scoring need to be implemented in private. Our code is publicly available  [^1] https://anonymous.4open.science/r/Private-Preserving_Data_Combination-451E.
@@ 3.4. SecureKL: Private KL-based Method
(Secure KL-X) $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}}=\mathbb{E}_{(x)\sim\mathcal{D}_{o}}(\text{Score}([x])).$
(Secure KL-XY) $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}=\mathbb{E}_{(x,y)\sim\mathcal{D}_{o}}(\text{Score}([x,y])).$
@@
(SecureKL Strategy, encrypted) $\pi_{p}(n=1)=\operatorname*{arg\,min}_{i\in[1..N]}\,g_{\mathrm{SKL}}(\mathcal{D}_{o},\mathcal{D}_{i}).$
@@
4. Experiments
4.1. Experimental Setup
We validate our method and demonstrate its applicability using the eICU Collaborative Research Dataset (Pollard2018TheEC, ), which contains over $200,000$ admissions from $208$ hospitals across the United States. Following the data cleaning and exclusion criteria outlined by (Water2023YetAI, ) and (shen2024data, ), we selected the $12$ hospitals with the highest number of patient visits (each with at least $2000$ patients) as our $\mathbf{H}$. Each strategy would compute with the same $K=3000$ records, as the total available data per hospital.
We simulate the problem setup for each hospital with the $24$-hour mortality prediction task. The strategy comparisons described in Section 3 are implemented using $1500$ samples and the AUC is evaluated on $400$ samples for all of our experiments unless otherwise noted. This follows training and evaluation protocols in Yet Another ICU Benchmark (Water2023YetAI, ). For the data combination experiments that compute AUC change $\delta_{i}$ or $\delta_{T}$, to match  (shen2024data, ) , we take $1500$ random samples from each selected dataset and combine it with $1500$ samples from $\mathcal{D}_{o}$ (fixed across all experiments).
Implementing $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ to be privacy-preserving requires training logistic regression model in private. This is used in the private setting to estimate $g_{\mathrm{SKL}}$$\mathrm{Score}([X])$ or $\mathrm{Score}([X,Y])$ – for each pair of hospitals. Our experiments train encrypted logistic regression in CrypTen (knott2021crypten, ) using the library’s SGD optimizer. To ensure a fair comparison between the scores obtained through plaintext and encrypted settings, we re-implement plaintext Score($X$) and Score($X,Y$) using logistic regression with SGD in PyTorch (paszke2019pytorch, ). This is because encrypted version of L-BFGS – the optimizer prior work (shen2024data, ) uses in plaintext-only with Scikit-Learn (scikit-learn, )– is not available in CrypTen, though it leads to better downstream performance. Hyper-parameter tuning for SGD in private and plain text are performed independently, with the details in Appendix B.
4.2. Experimental Questions

### v2

@@ Abstract
Access to diverse, high-quality datasets is crucial for machine learning model performance, yet data sharing remains limited by privacy concerns and competitive interests, particularly in regulated domains like healthcare. This dynamic especially disadvantages smaller organizations that lack resources to purchase data or negotiate favorable sharing agreements. We present SecureKL, a privacy-preserving framework that enables organizations to identify beneficial data partnerships without exposing sensitive information. Building on recent advances in dataset combination methods, we develop a secure multiparty computation protocol that maintains strong privacy guarantees while achieving $>90\%$ correlation with plaintext evaluations. In experiments with real-world hospital data, SecureKL successfully identifies beneficial data partnerships that improve model performance for intensive care unit mortality prediction while preserving data privacy. Our framework provides a practical solution for organizations seeking to leverage collective data resources while maintaining privacy and competitive advantages. These results demonstrate the potential for privacy-preserving data collaboration to advance machine learning applications in high-stakes domains while promoting more equitable access to data resources. Our code is publicly available at https://anonymous.4open.science/r/Private-Preserving_Data_Combination-451E.
1 Introduction
@@
Empirical scaling laws have established clear relationships between model performance and three key factors: compute, data, and model size [1, 2]. These relationships have driven remarkable improvements across computer vision, language processing, speech recognition, reinforcement learning, and healthcare applications [3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. Beyond simple scaling, the diversity of training data has proven crucial for enhancing model robustness to distribution shifts and mitigating performance disparities across demographic groups [13, 14].
@@
However, access to data varies significantly across entities and domains. While large tech companies have the data and compute resources to train the foundational models that now dominate general tasks, smaller players often lack such access. Domain-specific data is also becoming increasingly valuable for fine-tuning these general models [15, 16, 17], creating a competitive advantage for the data owners. As a result, entities with domain-specific data are more reluctant to share it for free, opting instead to sell it in emerging data markets [18, 19, 20]. This dynamic particularly disadvantages smaller organizations, which often lack both the resources to purchase data and the leverage to negotiate favorable sharing agreements.
@@
Scaling datasets in regulated domains is particularly challenging due to legal constraints and unpredictable outcomes from altering the training data composition. In healthcare, for example, patient data is heavily regulated by laws such as the Health Insurance Portability and Accountability Act (HIPAA), which imposes strict data-sharing constraints to protect patient privacy. Moreover, accumulating data from multiple sources introduces the risk of domain shift, where data from different distributions may degrade model performance instead of improving it [21, 22, 23]. Contrary to the intuition that more data always leads to better performance, combining datasets does not guarantee improvements — in fact, performance can decrease when incorporating additional data sources [24, 25, 13]. This non-monotonic behavior means that carefully selecting which datasets to combine is crucial, as using all available data may actually perform worse than using an optimal subset.
@@
While existing methods aim to identify datasets combinations that can improve performance, they often assume access to all relevant datasets [13, 26]. However, this is impractical because entities are reluctant to share their data due to privacy risks or the competitive value they associate with it. This reluctance to share creates a bottleneck in improving model performance. In high-stakes settings like healthcare, this bottleneck is particularly detrimental, as access to diverse data has the potential to drive significant advancements in patient care and outcomes [27, 28, 29, 30].
@@
  2. 2.
    We extend KL-XY score [13] with a secure multiparty protocol that enables privacy-preserving evaluation of potential data partnerships. Our protocol maintains strong privacy guarantees while utilizing the complete underlying datasets, achieving $>90\%$ correlation with plaintext computations.
  3. 3.
@@
  4. 4.
    In experimental scenarios requiring selection of three partner hospitals, Private-KL-XY outperforms alternative selection strategies including demographic-based selection (using gender, race, and age), and limited-sample plaintext selection.
@@ 2 Problem Setup
$\mathrm{AUC}_{o}=\max_{f(\theta)}\,\mathrm{AUC}(\mathcal{M},\mathcal{D}_{o})$ (Baseline Performance)
@@
$\mathrm{AUC}_{T}=\max_{f(\theta)}\,\mathrm{AUC}(\mathcal{M},\mathcal{D}_{T}).$ (Combined Performance)
@@
$\pi^{*}(\mathcal{D}_{o},n)=\operatorname*{arg\,max}_{T\subset{[1...N]\choose n}}\mathrm{AUC}_{T}$ (Ideal Dataset Combination)
@@
Our approach uses Kullback-Leibler (KL)-divergence-based methods to gauge data utility, building on prior work [13]. KL divergence [31], also called information gain [32], describes a measure of how much a model probability distribution $Q$ is different from a true probability distribution $P$:
$\mathrm{KL}(P||Q)=\int_{x\in\mathcal{X}}\log\frac{P(\mathop{}\!\mathrm{d}x)}{Q(\mathop{}\!\mathrm{d}x)}P(\mathop{}\!\mathrm{d}x)$ (Kullback–Leibler Divergence)
Because computing KL-divergence on datasets $\mathcal{D}_{o}$ and $\mathcal{D}_{i}$ is non-trivial,  [13] proposes two groups of scores to make this divergence approximation tractable from small samples. Specifically, score $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ first trains a logistic regression model on $\mathcal{D}_{o}\cup\mathcal{D}_{i}$ – where the labels are folded into the covariates — with the goal of inferring dataset membership. Then, the resulting model’s probability score function $\text{Score}(\cdot):\mathcal{X,Y}\to[0,1]$ is averaged over a dataset in $H_{o}$, obtaining
$\mathrm{KL}_{\mathcal{X}\mathcal{Y}}=\mathbb{E}_{(x,y)\sim\mathcal{D}_{o}}(\text{Score}(x,y)).$ (KL-XY Score)
@@
We operate under a semi-honest privacy model—also known as honest-but-curious or passive security—where parties follow protocols but may probe intermediate values. Parties are “curious”, meaning that they can probe into the intermediate values to avoid paying for the data. This assumes a weaker security model than malicious security where a corrupted party may input foul data, but ensures the algorithm to be private throughout the computation. This privacy preservation model incentivizes collaboration, improving upon methods in  [13] .
@@
To secure this divergence computation cryptographically, Secure Multiparty Computation (MPC) [33, 34] protocols are leveraged. Specifically, in $\mathrm{SecureKL}$, each party encodes $\mathcal{D}_{o}$ and $\mathcal{D}_{i}$ to preserve privacy for both parties. This is implemented with the research framework CrypTen [35], specialized for MPC and machine learning. Our algorithmic and engineering details are in Sections  3 and  4, respectively. For related secure techniques, see Section 5.4.
@@
3 Methods
Figure 1: Our Method $\mathrm{SecureKL}(\mathcal{D}_{o},\mathcal{D}_{i})$. Each side encrypts their data, and a model is trained on their joint data. Using private KL-based measures, their distance is computed, and the final result is revealed after decryption, which requires both parties.
@@ 3 Methods
Category 1, medium-to-high leakage, sharing raw data. $\pi_{s}(n,k)$ supposes each hospital to share a dataset of size $k$; a default setting of $1\%$ is commonplace practice in some contracts, as a pre-requisite to being considered [36]. Though leakage can be controlled through $k$, the data is inherently sensitive. The underlying distance uses  [13] ’s KL-XY Score.
@@
3.1 Trivial Baseline: Blind Selection
Blind selection refers to the process when no information is provided. $\pi_{0}(n)$ randomly selects $n$ disjoint data sources, until data purchasing budget runs out. This random strategy may evade selection biases and help gather diverse data. Yet, prior work  [13] suggests that $\pi_{0}(1)$ – randomly selecting one source – for ICU is risky and inefficient for mortality prediction.
3.2 Sharing Summary Statistics
@@
$\pi_{d}(n=1)=\operatorname*{arg\,min}_{i\in[1..N]}\,L_{2}(\phi(\mathcal{D}_{o})||\phi(\mathcal{D}_{i})).$ (Demographic-based Strategy)
3.3 KL-based Methods, in Plaintext
@@
This is implemented with KL-based measures similar to [13] . To recap, $\mathrm{KL}(P||Q)$ is not symmetrical, meaning that it is not a ”metric” that satisfies triangle inequality. Intuitively, this means the measure is directional: a hospital’s distribution $P_{o}$ may be ”close” to the target distribution $P_{i}$, but not the other way around:
$\mathrm{KL}(P_{o}||P_{i})=\int_{x\in\mathcal{X}}\log\frac{P_{o}(\mathop{}\!\mathrm{d}x)}{P_{i}(\mathop{}\!\mathrm{d}x)}P_{i}(\mathop{}\!\mathrm{d}x)$ (Ideal Estimator)
@@
Specifically, $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ score used in $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ first trains a logistic regression model [37] on $\mathcal{D}_{o}\cup\mathcal{D}_{i}$ – where the labels are folded into the covaraiates – with the goal of inferring dataset membership. A score of $0.5$ or less means the datasets are not distinguishable, making the data potentially useful.  [13] established the insight that in data-limited domains of heterogeneous data sources, domain shifts of the covariates are useful for predicting whether the additional data helps the original task. We note again that even though this model is trained on both parties’ data, the final algorithm that the hospital uses to train on combined data is not restricted.
@@
$\mathrm{KL}_{\mathcal{X}}=\mathbb{E}_{(x)\sim\mathcal{D}_{o}}(\text{Score}(x)).$ (KL-X)
$\mathrm{KL}_{\mathcal{X}\mathcal{Y}}=\mathbb{E}_{(x,y)\sim\mathcal{D}_{o}}(\text{Score}(x,y)).$ (KL-XY)
We focus on $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$, and reproduce [13] ’s results that it is predictive of downstream change in AUC.
@@
$\pi_{s}(n=1,k=K)=\operatorname*{arg\,min}_{i\in[1..N]}\,g_{\mathrm{KL}}(\mathcal{D}_{o},\mathcal{D}_{i}).$ (KL-based Strategy, in plaintext)
@@
3.4 SecureKL: Private KL-based Method
Using MPC, we extend on $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ to require no information sharing (besides what was already assumed public). Specifically we leverage the MPC based framework provided by CrypTen [35], a library designed for privacy-preserving machine learning, to implement private $\mathrm{KL}_{\mathcal{X}}$ and $\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$. As illustrated in Figure 1, the logistic regression as well as the scoring need to be implemented in private. Our code is publicly available  [^1] https://anonymous.4open.science/r/Private-Preserving_Data_Combination-451E.
@@ 3.4 SecureKL: Private KL-based Method
$\mathrm{Secure}\mathrm{KL}_{\mathcal{X}}=\mathbb{E}_{(x)\sim\mathcal{D}_{o}}(\text{Score}([x])).$ (Secure KL-X)
$\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}=\mathbb{E}_{(x,y)\sim\mathcal{D}_{o}}(\text{Score}([x,y])).$ (Secure KL-XY)
@@
$\pi_{p}(n=1)=\operatorname*{arg\,min}_{i\in[1..N]}\,g_{\mathrm{SKL}}(\mathcal{D}_{o},\mathcal{D}_{i}).$ (SecureKL Strategy, encrypted)
@@
4 Experiments
4.1 Experimental Setup
We validate our method and demonstrate its applicability using the eICU Collaborative Research Dataset [38], which contains over $200,000$ admissions from $208$ hospitals across the United States. Following the data cleaning and exclusion criteria outlined by [39] and [13], we selected the $12$ hospitals with the highest number of patient visits (each with at least $2000$ patients) as our $\mathbf{H}$. Each strategy would compute with the same $K=3000$ records, as the total available data per hospital.
We simulate the problem setup for each hospital with the $24$-hour mortality prediction task. The strategy comparisons described in Section 3 are implemented using $1500$ samples and the AUC is evaluated on $400$ samples for all of our experiments unless otherwise noted. This follows training and evaluation protocols in Yet Another ICU Benchmark [39]. For the data combination experiments that compute AUC change $\delta_{i}$ or $\delta_{T}$, to match  [13] , we take $1500$ random samples from each selected dataset and combine it with $1500$ samples from $\mathcal{D}_{o}$ (fixed across all experiments).
Implementing $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ to be privacy-preserving requires training logistic regression model in private. This is used in the private setting to estimate $g_{\mathrm{SKL}}$$\mathrm{Score}([X])$ or $\mathrm{Score}([X,Y])$ – for each pair of hospitals. Our experiments train encrypted logistic regression in CrypTen [35] using the library’s SGD optimizer. To ensure a fair comparison between the scores obtained through plaintext and encrypted settings, we re-implement plaintext Score($X$) and Score($X,Y$) using logistic regression with SGD in PyTorch [40]. This is because encrypted version of L-BFGS – the optimizer prior work [13] uses in plaintext-only with Scikit-Learn [41]– is not available in CrypTen, though it leads to better downstream performance. Hyper-parameter tuning for SGD in private and plain text are performed independently, with the details in Appendix 8.
4.2 Experimental Questions